OpenStudy (anonymous):

How many square inches is 15.75 feet?

OpenStudy (anonymous):

\[ \begin{array}l\color{red}{\text{ }}\color{orange}{\text{p}}\color{#E6E600}{\text{l}}\color{green}{\text{e}}\color{blue}{\text{a}}\color{purple}{\text{s}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{h}}\color{#E6E600}{\text{e}}\color{green}{\text{l}}\color{blue}{\text{p}}\color{purple}{\text{ }}\color{purple}{\text{m}}\color{red}{\text{e}}\color{orange}{\text{.}}\color{#E6E600}{\text{}}\end{array} \] @bbcream14

OpenStudy (calculusxy):

Do you have any choices?

OpenStudy (anonymous):

okay:)

OpenStudy (tkhunny):

\(15.75\;ft^{2}\cdot \left(\dfrac{12\;in}{1\;ft}\right)^{2}\) Multiply by the right form of one (1), and you have it.

OpenStudy (anonymous):

no. i am trying to solve a different question.......... here is the question.. What is the area in square inches of a rectangle with a length of 6.3 feet, and a width of 2.5 feet? A.189 in2 B.2268 in2 C.15.75 in2 D.141.75 in2

OpenStudy (anonymous):

1. The formula for finding the Area of a circle is Pie X R-squared. R 3 cm

OpenStudy (anonymous):

so 2268in²?

OpenStudy (anonymous):

squared is 9cm X 3,14 (value of Pie) = 28.26 square inches.2. The formula for

OpenStudy (anonymous):

finding the Area of a rectangle is Length X Width (LXW): 14 X 2 = 28 square inches.3.The answer is...900 ins. are in 75 ft. 24

OpenStudy (calculusxy):

But for your second question: Area of rectangle equals to length times width Multiply the numbers int feet for now. After you get the product, then you can multiply by 12 to convert to inches.

OpenStudy (anonymous):

thx @bbcream14

OpenStudy (anonymous):

and @calculusxy

OpenStudy (anonymous):

welcome sis:)(:

OpenStudy (calculusxy):

You are Welcome

OpenStudy (anonymous):

Simplify the expression. A. –20 B. –18 C. –8 D. 28

OpenStudy (anonymous):

wat expression

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):